OpenStudy (anonymous):

lim(x->inf) (((sqrt(n)((sqrt(3n+1) - sqrt(3n))) I keep getting DNE, but that is not the correct answer. any help is appreciated! Thanks.

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} [\sqrt{n}(\sqrt{3n+1}-\sqrt{3n})]\]

OpenStudy (anonymous):



OpenStudy (anonymous):

@helpme1.2 i wrote it again in equation, and the sqrt(n) is multiplied by (sqrt(3n+1)-sqrt(3n))

OpenStudy (anonymous):

\[(\sqrt{n})*(\sqrt{3n+1}-\sqrt{3n})\]